OpenStudy (anonymous):

What is the remainder when 15,908 is divided by 13?

OpenStudy (plasmataco):

well, divide the number

OpenStudy (anonymous):

i get 28.32552

OpenStudy (plasmataco):

without a calculator

OpenStudy (anonymous):

ok

TheSmartOne (thesmartone):

Here's a tip. You do it with a calculator. What do you get when you do it with a calculator? No, it is not 28.32552

OpenStudy (plasmataco):

and the decimal version to this problem is 1223.6923

OpenStudy (plasmataco):

well?

OpenStudy (anonymous):

Yea i accidentally looked at something else i realized a couple mins ago

OpenStudy (plasmataco):

so do u know the answer?

OpenStudy (anonymous):

Nope .

OpenStudy (anonymous):

ohh oh myyy god

OpenStudy (plasmataco):

wat?

OpenStudy (anonymous):

9 xD

OpenStudy (plasmataco):

Nice work!

OpenStudy (anonymous):

:) yas

OpenStudy (anonymous):

tyty very mucho

OpenStudy (plasmataco):

no prob

OpenStudy (anonymous):

*face palm* im so slow i swear

OpenStudy (plasmataco):

I'm am a very good teacher

OpenStudy (plasmataco):

XD

OpenStudy (anonymous):

xD Yeah

Zeronknight (zeronknight):

\(\Huge\cal\color{lightgreen}{Welcome~}\color{red}{To~}\color{lime}{OpenStudy!}\)

OpenStudy (plasmataco):

yep, where people help thoroughly teach the problems you have... or else the moderators kik em off

OpenStudy (anonymous):

rude ;O

OpenStudy (plasmataco):

lel yep

TheSmartOne (thesmartone):

so we got 1223.6923 now 15908/ 13 = 1223.6923 now, we basically do 13 * 1223 = ?? What is 13 * 1223 = ?

OpenStudy (plasmataco):

lel, yjgg became a master the old fasion way XD

jimthompson5910 (jim_thompson5910):

Alternative way 15908/13 = 1,223.6923076923 focus on the decimal portion 0.6923076923 multiply it by 13 to get 0.6923076923*13 = 8.9999999999 which essentially rounds to 9 It would be exactly 9 if we didn't have rounding error

OpenStudy (anonymous):

Yeah thank you everyone ya'll so helpful :)

OpenStudy (plasmataco):

no prob

TheSmartOne (thesmartone):

what I was going to say would be to 13 * 1223 and then whatever you get, you subtract it from 15908 15908 - (13*1223) = your remainder